OpenStudy (anonymous):

Rationalize the denominator: (3^1/2 +2) / ( b^1/4 +c )

OpenStudy (anonymous):

\[\sqrt{3}+2\div \sqrt{\sqrt{b}}+c\]

OpenStudy (anonymous):

 is this correct